OpenStudy (anonymous):

solve the equation t(4t+1)=0

OpenStudy (callisto):

t(4t+1)=0 t=0 or 4t+1 = 0 t=-1/4 therefore t = 0 or t=-1/4

OpenStudy (anonymous):

t=0 or t=-1/4

OpenStudy (anonymous):

t(4t+1)=0 then t=0 or 4t+1=0 therefore t=-1/4 t=0

OpenStudy (anonymous):

solution take T and (4t+1) = 0 s0 t = 0 4t+1=0 4t=-1 t=-1/4